A chemistry grad student measures the performance Z of the new pump in his lab. The result is:
Z=2.2 kPa*mm^3 * s^-1

Convert Z to mJ*s^-1
1
Expert's answer
2019-08-23T06:34:40-0400

"2.2 KPa \times mm^3 \times sec^{-1}"

"\to 2.2 \times 10^3Pa \times (10^{-3})^3 m^3 \times sec^{-1}"

"\to 2.2 \times 10^{-6}m^3 Pa \space sec^{-1}" "(m^3Pa=J)"

"\to 2.2 \times 10^{-6} J \space sec^{-1} (1 mJ=10^3 J)"

"\to 2.2 \times 10^{-6} \times 10^3 mJ sec^{-1}"

"\to 2.2\times 10^{-3} mJ \space sec^{-1}"